OpenStudy (anonymous):

Fool's problem of the day, Today, it's a cute number theory problem, Can you find the last digit \( 77{\uparrow\uparrow}9 \)? [Solved by @KingGeorge] Notation Reference: \( \large \begin{matrix} a\uparrow\uparrow b & = {\ ^{b}a} = & \underbrace{a^{a^{{}^{.\,^{.\,^{.\,^a}}}}}} & \\ & & b\mbox{ copies of }a \end{matrix} \) Good luck!

OpenStudy (asnaseer):

what does the double up arrow signify?

OpenStudy (kinggeorge):

\[\Large 77 \uparrow \uparrow 9=77^{77^{77^{...^{77}}}}\]Where there are 9 "77's" in the stack. Since we're looking mod for the last digit, we need to take mod 10.

OpenStudy (asnaseer):

knowing FFM I'm sure we need to make use of primes factors somewhere: 77 = 7 * 11

OpenStudy (asnaseer):

11 raised to any power will always have last digit equal to 1

OpenStudy (asnaseer):

so we just need to concentrate on the 7 I would guess

OpenStudy (anonymous):

@asnaseer: I love primes! But the general solution i have for any \( x{\uparrow\uparrow}a\) doesn't uses prime factorization.

OpenStudy (asnaseer):

FFM: you have a GENERAL solution for \(x{\uparrow\uparrow}a\) !!! you MUST be from another planet! :D

OpenStudy (kinggeorge):

This is equivalent to asking \[\Large 7^{7^{7^{...^7}}} \mod 10\] with 9 7's. Since \(7^4 \equiv 1 \mod 10\), we need to find \[7 \uparrow\uparrow8 \mod 4\]Then, we know that \(7 \equiv -1 \mod 4\), so we're finding \[\Large -1^{3^{3^{...^3}}} \mod 4\]Since 3 to any power is odd, this is equivalent to -1. This leaves us with \(7^{-1} \mod 10\) This is easy to calculate, and is equivalent to 3.

OpenStudy (anonymous):

@KingGeorge: What about the general problem?

OpenStudy (anonymous):

What level of math is this?

OpenStudy (kinggeorge):

In general, I'll have to think about it some more. I've got class now, so I'll return later hopefully with a solution. btw, I am correct in thinking 3 was the answer right?

OpenStudy (anonymous):

Yes you are.

OpenStudy (experimentx):

looks like out of my league

OpenStudy (anonymous):

i have no experience with modular arithmatic, am i allowed to have mod functions within other functions?

OpenStudy (kinggeorge):

Back now. While I'm not positive about getting a formula for \(a\uparrow\uparrow b \mod 10\) for all \(a, b\), I have noticed some striking patterns. First pattern (trivial):\[a \uparrow\uparrow 1\equiv a \mod 10\] Second pattern: If the \(\gcd(a, 10)=1\) and \(b \geq2\),\[a\uparrow\uparrow b\equiv a^{-1} \mod 10\] Pattern three: If \(a=5\) or \(a=6\)\[a\uparrow\uparrow b =a \mod 10\] When a equals the powers of 2 however, it gets a little funky. For 2, \[2 \uparrow\uparrow b=5+(-1)^{b-1} \mod 10 \qquad\quad b\geq2\]So it alternates between 4 and 6. When \(a=4=2^2\), \[4 \uparrow\uparrow b \equiv 5+(-1)^{2(b-1)}\equiv5+1\equiv6 \mod 10 \qquad\quad b\geq2\]Finally, if \(a=8=2^3\), we have that \[8 \uparrow\uparrow b= 5+(-1)^{3(b-1)} \equiv 5-(-1)^{b-1} \mod10\qquad\quad b\geq2\]So it alternates between 4 and 6 in the opposite manner of \(a=2\).

OpenStudy (kinggeorge):

It should also be noted that \(a\uparrow\uparrow b\) can be defined recursively as\[\Large a\uparrow\uparrow b=a^{a\uparrow\uparrow (b-1)}\]